A recent study found that 2 out of every 10 houses in a neighborhood have no insurance. If 5 houses are selected from 10 houses, find the probability that fewer than 4 will be insured.


1
Expert's answer
2022-01-18T13:59:44-0500

for binomial distribution we have:

"p=0.8,n=5"

"P(x=k)=C^k_n p^k (1-p)^{n-k}"


"P(x<4)=P(0)+P(1)+P(2)+P(3)"

"P(0)=0.2^5=0.0003"

"P(1)=5\cdot0.8\cdot0.2^4=0.0064"

"P(2)=10\cdot0.8^2\cdot0.2^3=0.0512"

"P(3)=10\cdot0.8^3\cdot0.2^2=0.2048"


"P(x<4)=0.0003+0.0064+0.0512+0.2048=0.2627"